OpenStudy (anonymous):

How do you solve this quadratic equation ? (2x-1)(3x-1) = 0 Can you please explain it to me...?

OpenStudy (a1234):

Is that 2x -1?

OpenStudy (anonymous):

1/2 and 1/3

OpenStudy (a1234):

Okay... add the 1s to the zero. 1 = 2x 1 = 3x Divide each side by three, and you get: x = 1/2 and x = 1/3.

OpenStudy (anonymous):

ohhhh, okay. thanks (: can you help me with this one too ? 3x^2 + 12x + 9 = 0

OpenStudy (a1234):

Find two numbers that have a product of 27 (9*3) and a sum of 12. 3x^2 + 9x + 3x + 9 = 0 Pull out the GCF: 3x(x + 3) + 3(x + 3) = 0 (3x + 3)(x + 3) = 0 x = -3 3x = -3 x = -1 x = -1 and x = -3.

OpenStudy (anonymous):

thankssss

OpenStudy (a1234):

You're welcome.